### v1

@@ Abstract
The task of medical image recognition is notably complicated by the presence of varied and multiple pathological indications, presenting a unique challenge in multi-label classification with unseen labels. This complexity underlines the need for computer-aided diagnosis methods employing multi-label zero-shot learning. Recent advancements in pre-trained vision-language models (VLMs) have showcased notable zero-shot classification abilities on medical images. However, these methods have limitations on leveraging extensive pre-trained knowledge from broader image datasets, and often depend on manual prompt construction by expert radiologists. By automating the process of prompt tuning, prompt learning techniques have emerged as an efficient way to adapt VLMs to downstream tasks. Yet, existing CoOp-based strategies fall short in performing class-specific prompts on unseen categories, limiting generalizability in fine-grained scenarios. To overcome these constraints, we introduce a novel prompt generation approach inspirited by text generation in natural language processing (NLP). Our method, named Pseudo-Prompt Generating (PsPG), capitalizes on the priori knowledge of multi-modal features. Featuring a RNN-based decoder, PsPG autoregressively generates class-tailored embedding vectors, i.e., pseudo-prompts. Comparative evaluations on various multi-label chest radiograph datasets affirm the superiority of our approach against leading medical vision-language and multi-label prompt learning methods. The source code is available at https://github.com/fallingnight/PsPG.
@@ 1 Introduction
Medical images like chest radiographs frequently contain multiple pathological features in real-world scenarios[6, 9, 27]. Moreover, specific contexts or annotation strategies can lead to open-vocabulary pathology categorizations [^1] Conditions such as “Pleural Thickening” and “Pleural Adhesions”, for instance, may be grouped into a broader “Pleural Other” category[9].. This complexity underlines the need for computer-aided diagnosis methods employing multi-label zero-shot learning to provide more reliable clinical decision-making support. Yet, the high costs associated with expert labeling of extensive chest radiograph datasets pose a barrier, resulting in a scarcity of large-scale, realistically annotated datasets. Despite these challenges, large medical datasets like MIMIC-CXR[11] offer a wealth of unstructured medical report texts, enabling the exploration of self-supervised image-text contrastive learning methods[8, 17, 23, 34]. These methods develop a visual-linguistic associative latent space through extensive pre-training. During inference, they capitalize on the similarities between the textual features of classificatory prompts and the visual features of images, facilitating classification to extend to unseen categories.
@@ 3.2 Pseudo-Prompt Generating
Figure 2: Overview of the prompt learning phase of PsPG.
@@
The PsPG training unfolds in two phases. Initially, it follows CLIP’s pre-training to perform fine-tuning on a medical image-text dataset as in 3.1. The second phase, central to our method, involves prompt learning on an image-label dataset, as depicted in Fig. 2. Prompt Decoder generates class-tailored pseudo-prompts from visual global features and class textual features, the latter illustrated in Fig. 1 (c). Note that what we generate is not prompt templates, but complete pseudo-prompts, which do not need to be concatenated with the corresponding class names. Pseudo-prompts directly feed into the text encoder that bypasses the first embedding layer. Two prompt decoders operate in parallel to derive both positive and negative pseudo-prompts.
@@
where $\mathrm{\Omega}\in\mathbb{Z}^{N_{c}\times N_{c}}$. $\mathrm{\Omega}$ is a symmetric matrix, and the element $\omega_{ij}$ of $\mathrm{\Omega}(i\in[1,Nc],j\in[1,Nc],i<j)$ denotes the times of co-occurrences of label $l_{i}$ and $l_{j}$ within the mini-batch. For simplicity, we construct correlation targets $Q$ by flattening the strictly upper triangular matrix of $\mathrm{\Omega}$ and setting all $\omega_{ij}>1$ to 1, $Q\in\{0,1\}^{C_{2}^{N_{c}}}$.
Fig. 2 illustrates that using the pairwise concatenation of pseudo-prompts $P_{i}P_{j}$ (where $i,j\in[1,N_{c}]$ and $i<j$) to yield $W_{pair}$, with $W_{pair}\in\mathbb{R}^{C_{2}^{N_{c}}\times D_{vlp}}$. For $X_{u}$ and its global-local feature $U_{gl}$, the probability $p_{mn}$ is given by:
@@
We define a batch-level co-occurrence target $Q$, serving as the soft target for all label pair co-occurrence probabilities within the mini-batch. Maximum pooling is applied to these probabilities to computing the Binary Cross-Entropy Loss with the soft target $Q$. The overall loss $L$ for PsPG is:
@@ 4.1 Setup
Private-CXR is a private chest radiograph dataset collected from The First People’s Hospital of Hangzhou Lin’an District, containing 7,357 images and 18 pathology categories all labelled by experienced radiologists. Contingent on participant consent, we used the entire dataset for zero-shot evaluation.
@@
Medical Vision-Language Pre-training Methods: These pioneering contrastive learning methods, include ConVIRT[34], GLoRIA[8], BioViL-T[4], MedCLIP[28], Chexzero[23] and KAD[32], were known for their efficacy in zero-shot medical image classification [^2] We extended the originally single-label focused to multi-label, as detailed in 3.1.. Pre-trained weights were utilized where available. For which absent such weights, we retrained it on MIMIC-CXR. Additionally, we compared with the fine-tuned CLIP prior to prompt tuning.
@@
Aligning with most baselines, we used ResNet50 backbone weights provided by CLIP[17]. In backbone fine-tuning, we use AdamW optimizer with weight decay 0.05, and set learning rate to $1\times 10^{-4}$ with cosine annealing. We trained 15 epochs, with batch size 64, 5 warm-up epochs and warm-up learning rate of $1\times 10^{-6}$. With a frozen backbone, we tuned the prompt decoder (sequence length 16, hidden dimension 512) and set ASL $\gamma^{+}=1$, $\gamma^{-}=4$, and $c=0.05$. We use SGD optimizer, and set learning rate to $1\times 10^{-4}$ with cosine annealing. We trained 50 epochs, with batch size 64, 5 warm-up epochs and warm-up learning rate of $1\times 10^{-6}$. We select the best epoch by evaluating the Macro AUC performance on the CheXpert-val. All training phases are done with one NVIDIA RTX 3080Ti.
@@
For PsPG, KAD, and DualCoOp, we used class names as input. For other baselines, we applied fixed templates from BioViL [^3] Positive: “Findings suggesting {cls}.” Negative: “No evidence of {cls}.”. Similar to typical medical tasks and multi-label tasks, we use Macro AUC, Micro AUC, mAP as evaluation metrics. Predictions confidence intervals were computed using a non-parametric bootstrap with 1000 re-samples and $\alpha$ of 0.05. Due to page limitations, we placed the detailed confidence intervals in the supplementary material.
@@ 4.2 Multi-label Zero-shot Classification
To make evaluation categories consistent, we evaluated Generalized Zero-Shot Learning (GZSL) on four datasets. Unlike the CheXpert original study’s focus on five competition categories[9], Our evaluation spans all 14 pathology categories. In Table 1, $\text{CLIP}_{\text{fine-tuned}}$ denotes the CLIP we fine-tune on the medical dataset. $\text{DualCoOp}_{\text{init}}$ denotes template initialization using 33footnotemark: 3, and $\text{DualCoOp}_{\text{16}}$ denotes using the same length of prompt sequences as ours. $\textbf{PsPG}_{\text{prefix}}$ denotes using our method to learn prefixes than whole prompts. Additionally, we compare the performance of fine-tune the pre-trained CheXpert weights on ChestXray-14[27] for $\text{DualCoOp}_{\text{cross}}$ and $\textbf{PsPG}_{\text{cross}}$. More detailed results about the cross-dataset generalization are shown in supplementary material.
Table 1: Multi-label Zero-shot Classification on CheXpert, VinDr-CXR, Private-CXR, PadChest. #P stands for trainable parameters of the module. All metrics are displayed in percentage. For PadChest, we only evaluate on categories which positive samples $n_{pos}>50$ (108 categories). We underlined the best performance across all methods (except cross methods).All models except BioViL-T* (only provided $512\times 512$) evaluated used $224\times 224$ as visual input.

### v2

@@ Abstract
The task of medical image recognition is notably complicated by the presence of varied and multiple pathological indications, presenting a unique challenge in multi-label classification with unseen labels. This complexity underlines the need for computer-aided diagnosis methods employing multi-label zero-shot learning. Recent advancements in pre-trained vision-language models (VLMs) have showcased notable zero-shot classification abilities on medical images. However, these methods have limitations on leveraging extensive pre-trained knowledge from broader datasets, and often depend on manual prompt construction by expert radiologists. By automating the process of prompt tuning, prompt learning techniques have emerged as an efficient way to adapt VLMs to downstream tasks. Yet, existing CoOp-based strategies fall short in performing class-specific prompts on unseen categories, limiting generalizability in fine-grained scenarios. To overcome these constraints, we introduce a novel prompt generation approach inspirited by text generation in natural language processing (NLP). Our method, named Pseudo-Prompt Generating (PsPG), capitalizes on the priori knowledge of multi-modal features. Featuring a RNN-based decoder, PsPG autoregressively generates class-tailored embedding vectors, i.e., pseudo-prompts. Comparative evaluations on various multi-label chest radiograph datasets affirm the superiority of our approach against leading medical vision-language and multi-label prompt learning methods. The source code is available at https://github.com/fallingnight/PsPG.
@@ 1 Introduction
Medical images like chest radiographs frequently contain multiple pathological features in real-world scenarios[6, 9, 27]. Moreover, specific contexts or annotation strategies can lead to open-vocabulary pathology categorizations [^1] Conditions such as “Pleural Thickening” and “Pleural Adhesions”, for instance, may be grouped into a broader “Pleural Other” category[9].. This complexity underlines the need for computer-aided diagnosis methods employing multi-label zero-shot learning to provide more reliable clinical decision-making support. Yet, the high costs associated with expert labeling of extensive chest radiograph datasets pose a barrier, resulting in a scarcity of large-scale, realistically annotated datasets. Despite these challenges, large medical datasets like MIMIC-CXR[11] offer a wealth of unstructured medical report texts, enabling the exploration of self-supervised image-text contrastive learning methods[8, 17, 23, 34]. These methods develop a visual-linguistic associative latent space through extensive pre-training. During inference, they capitalize on the similarities between features of classificatory prompts and images, facilitating classification to extend to unseen categories.
@@ 3.2 Pseudo-Prompt Generating
Figure 2: Overview of the prompt learning phase of PsPG. We introduce three novel components: Prompt Decoder, Spatial Fusion, and Soft Pairwise Co-occurrence Loss.
@@
The PsPG training unfolds in two phases. Initially, it follows CLIP’s pre-training to perform fine-tuning on a medical image-text dataset as in 3.1. The second phase, central to our method, involves prompt learning on an image-label dataset, as depicted in Fig. 2. Prompt Decoder generates class-tailored pseudo-prompts from visual global features and class textual features, illustrated in Fig. 1 (c). Note that what we generate is not prompt templates, but complete pseudo-prompts, which do not need to be concatenated with the corresponding class names. Pseudo-prompts directly feed into the text encoder that bypasses the first embedding layer. Two prompt decoders operate in parallel to derive both positive and negative pseudo-prompts.
@@
where $\mathrm{\Omega}\in\mathbb{Z}^{N_{c}\times N_{c}}$. $\mathrm{\Omega}$ is a symmetric matrix, and the element $\omega_{ij}$ of $\mathrm{\Omega}(i\in[1,Nc],j\in[1,Nc],i<j)$ denotes the times of co-occurrences of label $l_{i}$ and $l_{j}$ within the mini-batch. For simplicity, we construct co-occurrence targets $Q$ by flattening the strictly upper triangular matrix of $\mathrm{\Omega}$ and setting all $\omega_{ij}>1$ to 1, $Q\in\{0,1\}^{C_{2}^{N_{c}}}$.
Fig. 2 illustrates that using the pairwise concatenation of pseudo-prompts $P_{i}P_{j}$ (where $i,j\in[1,N_{c}]$ and $i<j$) to yield $W_{pair}$, with $W_{pair}\in\mathbb{R}^{C_{2}^{N_{c}}\times D_{vlp}}$. For the global-local feature $U_{gl}$ of a batch, the probability $p_{mn}$ is given by:
@@
We use the batch-level co-occurrence target $Q$ to serve as the soft target for all label pair co-occurrence probabilities within the mini-batch. Maximum pooling is applied to these probabilities to computing the Binary Cross-Entropy Loss with the soft target $Q$. The overall loss $L$ for PsPG is:
@@ 4.1 Setup
Private-CXR is a chest radiograph dataset collected from The First People’s Hospital of Hangzhou Lin’an District, containing 7,357 images and 18 pathology categories all labelled by experienced radiologists. The availability of this dataset is restricted. Contingent on participant consent, we used the entire dataset for zero-shot evaluation.
@@
Medical Vision-Language Pre-training Methods: These pioneering contrastive learning methods, include ConVIRT[34], GLoRIA[8], BioViL-T[4], MedCLIP[28], Chexzero[23] and KAD[32], were known for their efficacy in zero-shot medical image classification [^2] We extended the originally single-label focused to multi-label, as detailed in 3.1.. Pre-trained weights were utilized where available. For which absent such weights[34], we retrained it on MIMIC-CXR. Additionally, we compared with the fine-tuned CLIP prior to prompt tuning.
@@
Aligning with most baselines, we used ResNet50 backbone weights provided by CLIP[17]. In backbone fine-tuning, we use AdamW optimizer with weight decay 0.05, and set learning rate to $1\times 10^{-4}$ with cosine annealing. We trained 15 epochs, with batch size 64, 5 warm-up epochs and warm-up learning rate of $1\times 10^{-6}$. With a frozen backbone, we tuned the prompt decoder (sequence length 16, hidden dimension 512) and set ASL $\gamma^{+}=1$, $\gamma^{-}=2$, and $c=0.05$. We use SGD optimizer, and set learning rate to $1\times 10^{-4}$ with cosine annealing. We trained 50 epochs, with batch size 64, 5 warm-up epochs and warm-up learning rate of $1\times 10^{-6}$. We select the best epoch by evaluating the Macro AUC performance on the CheXpert-val. All training phases are done with one NVIDIA RTX 3080Ti.
@@
For PsPG, KAD, and DualCoOp, we used class names as input. For other baselines, we applied fixed templates from BioViL [^3] Positive: “Findings suggesting {cls}.” Negative: “No evidence of {cls}.”. Similar to typical medical tasks and multi-label tasks, we use Macro AUC, Micro AUC, mAP as evaluation metrics. Predictions confidence intervals were computed using a non-parametric bootstrap with 1000 re-samples and $\alpha$ of 0.05. Due to page limitations, we placed the detailed confidence intervals in appendix.
@@ 4.2 Multi-label Zero-shot Classification
To make evaluation categories consistent, we evaluated Generalized Zero-Shot Learning (GZSL) on four datasets. Unlike the CheXpert original study’s focus on five competition categories[9], Our evaluation spans all 14 pathology categories. In Table 1, $\text{CLIP}_{\text{fine-tuned}}$ denotes the CLIP we fine-tune on the medical dataset. $\text{DualCoOp}_{\text{init}}$ denotes template initialization using 33footnotemark: 3, and $\text{DualCoOp}_{\text{16}}$ denotes using the same length of prompt sequences as ours. $\textbf{PsPG}_{\text{prefix}}$ denotes using our method to learn prefixes than whole prompts. Additionally, we compare the performance of fine-tune the pre-trained CheXpert weights on ChestXray-14[27] for $\text{DualCoOp}_{\text{cross}}$ and $\textbf{PsPG}_{\text{cross}}$.
Table 1: Multi-label Zero-shot Classification on CheXpert, VinDr-CXR, Private-CXR, PadChest. #P stands for trainable parameters of modules. All metrics are displayed in percentage. For PadChest, we only evaluate on categories which positive samples $n_{pos}>50$ (108 categories). We underlined the best performance across all methods (except cross methods).All models except BioViL-T* (only provided $512\times 512$) evaluated used $224\times 224$ as visual input.